OpenStudy (anonymous):

What is the vertex of f(x) = x2 + 10x + 16? (Points : 2) (2, 6) (1, 5) (5, 9) (-5, -9)

OpenStudy (anonymous):

don't suppose you can take the derivative and set it to zero?

OpenStudy (turingtest):

I doubt it, think they're supposed to complete the square

OpenStudy (turingtest):

..or use one of those formulas

OpenStudy (mathmate):

If you do completing the squares, you'll get it into the form y=a(x-h)^2+k where (h,k) is the vertex. Since h=-5, there's only one choice that matches. x^2+10x+...=(x+5)^2+...., so h=-5

OpenStudy (anonymous):

Thankss mathmate you made mathkid very happy :)

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (across):

There's really no need to complete the squares. :P\[x^2+10x+16=(x+2)(x+8)\]\(x=-2\) or \(x=-8\).\[\frac{(-8)+(-2)}{2}=-5!\]